A bird is sitting in a tree 4m in the air. It has 220 J energy. How massive is the bird?


1
Expert's answer
2021-03-14T19:14:04-0400
"PE=mgh,""m=\dfrac{PE}{gh}=\dfrac{220\ J}{9.8\ \dfrac{m}{s^2}\cdot4\ m}=5.61\ kg."